# Quantifying the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data

### v1

@@ Appendix B Preliminary Experiment Results
The K-S test is a non-parametric statistical test used to compare the distribution of two datasets. It evaluate whether two samples come from the same underlying probability distribution. The null hypothesis of the K-S test is that the two samples are drawn from the same distribution, while the alternative hypothesis is that the two samples are drawn from different distributions. There are two key values provided by K-S test: the K-S Statistic quantifies the largest difference between the two sample distributions, and the p-value assess the statistical significance of that difference. If the p-value is lower than a chosen significance level (0.05 in our experiments), we reject the null hypothesis, concluding that the two distributions are significantly different. Otherwise, we fail to reject the null hypothesis, suggesting that there is no significant difference between the two distributions. We provide the K-S test results in Table 3 and Table 4, with the corresponding distribution visualizations plots shown in Figure 6 and Figure 7.
@@
[…]
@@
Appendix C Implementation Details for Injecting Spurious Features
We provide detailed prompts for LLM-based perturbations in Figure 8. For rule-based perturbations, placeholder template is presented in Figure 9.
@@ Appendix C Implementation Details for Injecting Spurious Features
Appendix D Implementation Details for Preserving Causal Features
We employ a bidirectional entailment algorithm to ensure the semantic equivalence before and after introducing spurious features. The prompts for its core component, NLI model, are shown in Figure 10.
@@ Appendix D Implementation Details for Preserving Causal Features
Appendix E Statistics of the Synthetic Dataset
@@
Table 5: Distribution of the synthetic dataset for Llama-3.1-8B-Instruct.
Appendix F Experimental Setup Details
Dataset
We conduct our main experiments using the synthetic dataset generated by the SURE framework. The queries are sourced from the NQ-open dataset  [^6] https://huggingface.co/datasets/google-research-datasets/nq_open, and the documents are derive from the English Wikipedia dump.
Models
We test two representative LLMs in our main experiments: Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct.
@@ Appendix F Experimental Setup Details
The instruction $I$ in the RAG prompt $P=(I,G,Q)$, shown in Figure 11, is derived from Cuconasu et al. (2024), with slight modifications to better adapt to our setting.
@@
We follow the typical ”retrieve-read” setting of RAG paradigm. For the retrieval module, we use Contriever-msmarco [^7] https://huggingface.co/facebook/contriever-msmarco, a BERT-based dense retriever, as the default retriever. It is finetuned on the MS MARCO dataset (Bajaj et al., 2016) after unsupervised pretraining via contrastive learning (Izacard et al., 2021). To optimize the efficiency of vector similarity searches, we employ the Faiss library (Douze et al., 2024). For the read module, we deploy LLMs on NVIDIA A100 GPUs and accelerate inference with vllm [^8] https://github.com/vllm-project/vllm. We follow (Li et al., 2023b; Zhang et al., 2024a, b) in using a greedy decoding strategy and set the temperature to 0.1 to ensure stable outputs and strong reproducibility.
@@
Figure 11: Instruction $I$ used for the QA task.
@@
Spurious Features Perturbations Known-Golden Known-Noise Unknown-Golden Unknown-Noise Total
LR RR WR LR RR WR LR RR WR LR RR WR LR RR WR
Style Simple 7.79 83.04 9.18 1.70 95.80 2.50 8.43 82.88 8.69 0.27 99.45 0.28 1.80 95.96 2.24
Complex 6.00 85.60 8.40 1.91 96.59 1.50 6.71 84.86 8.43 0.23 99.57 0.20 1.63 96.62 1.75
Source LLM-Generated 5.89 86.43 7.69 1.43 96.83 1.74 6.20 85.71 8.09 0.22 99.56 0.22 1.51 96.60 1.89
Self-Generated 6.55 85.01 8.44 1.55 96.37 2.09 6.52 86.36 7.12 0.19 99.57 0.24 1.62 96.32 2.06
Logic Reverse 5.06 90.82 4.12 1.13 97.82 1.06 5.73 89.71 4.56 0.21 99.67 0.13 1.29 97.64 1.07
Random 3.91 93.16 2.93 0.86 98.31 0.83 4.21 91.67 4.12 0.14 99.72 0.14 0.97 98.18 0.85
LLM-Ranked 3.24 93.93 2.83 0.82 98.43 0.74 3.58 93.36 3.06 0.13 99.76 0.11 0.85 98.39 0.75
Format JSON 7.01 88.25 4.74 1.70 97.25 1.05 5.92 89.63 4.45 0.25 99.61 0.14 1.76 97.08 1.16
HTML 11.85 84.46 3.69 2.70 96.90 0.40 9.33 86.78 3.90 0.35 99.61 0.04 2.85 96.41 0.74
YAML 5.26 89.94 4.80 1.26 97.41 1.33 4.79 90.80 4.41 0.17 99.67 0.16 1.32 97.40 1.28
Markdown 2.32 92.23 5.45 0.60 96.89 2.51 2.34 93.46 4.19 0.07 99.61 0.32 0.61 97.55 1.84
Metadata Timestamp (pre) 2.08 95.81 2.11 0.28 99.42 0.29 2.54 95.56 1.90 0.02 99.95 0.03 0.46 99.10 0.44
Timestamp (post) 2.04 95.86 2.10 0.25 99.43 0.32 2.81 95.56 1.63 0.02 99.95 0.03 0.46 99.11 0.43
Datasource (wiki) 2.11 93.45 4.44 0.23 98.96 0.81 3.25 92.47 4.27 0.03 99.86 0.11 0.48 98.50 1.03
Datasource (twitter) 2.27 94.11 3.62 0.31 99.25 0.43 2.77 93.97 3.25 0.02 99.91 0.07 0.50 98.77 0.73

### v2

@@ A.3 Preliminary Experiment Results
The K-S test is a non-parametric statistical test used to compare the distribution of two datasets. It evaluate whether two samples come from the same underlying probability distribution. The null hypothesis of the K-S test is that the two samples are drawn from the same distribution, while the alternative hypothesis is that the two samples are drawn from different distributions. There are two key values provided by K-S test: the K-S Statistic quantifies the largest difference between the two sample distributions, and the p-value assess the statistical significance of that difference. If the p-value is lower than a chosen significance level (0.05 in our experiments), we reject the null hypothesis, concluding that the two distributions are significantly different. Otherwise, we fail to reject the null hypothesis, suggesting that there is no significant difference between the two distributions.
The K-S statistic and P-value are presented in Table 6 and Table 7. Furthermore, we visualize the feature distributions for both the experimental and control groups in Figure 6. For all tested features in the experimental group, the K-S test rejects the null hypothesis, concluding that the distribution of the two sets are significantly different. In contrast, for the control group, the K-S test fails to reject the null hypothesis. The results for Llama-3.1-8B-Instruct are also provided in Figure 7. According to these results, we can conclude that RALMs exhibit bias toward spurious features in documents.
@@
Flesch score 0.0677 $1.01\text{\times}{10}^{-5}$∗∗∗ 0.0301 0.1799
Distinct-1 0.0756 $4.95\text{\times}{10}^{-7}$∗∗∗ 0.0203 0.6431
DTD 0.0636 $4.29\text{\times}{10}^{-5}$∗∗∗ 0.0124 0.9866
PPL 0.0722 $1.88\text{\times}{10}^{-6}$∗∗∗ 0.0162 0.8776
Token Length 0.1708 $2.91\text{\times}{10}^{-34}$∗∗∗ 0.0256 0.3493
@@
Appendix B Implementation Details for Injecting Spurious Features
We provide detailed prompts for LLM-based perturbations in Figure 8. For rule-based perturbations, placeholder templates are presented in Figure 9.
@@ Appendix B Implementation Details for Injecting Spurious Features
Appendix C Implementation Details for Preserving Causal Features
We employ a bidirectional entailment algorithm to ensure the semantic equivalence before and after introducing spurious features. The prompts for its core component, NLI model, are shown in Figure 10. Furthermore, we apply a simple string-matching strategy to preserve ground truths. Specifically, for Golden documents that originally contained the correct answers, we keep them only if they preserve the ground truths after perturbation. For Noise documents that initially lack the correct answers, we discard them if they unexpectedly acquire ground truths due to perturbations.
@@ Appendix C Implementation Details for Preserving Causal Features
Appendix D Statistics of the Synthetic Dataset
@@
Table 8: Distribution of the SURE_Wiki dataset for Llama-3.1-8B-Instruct.
Appendix E Experimental Setup Details
@@
The instruction II in the RAG prompt $P=(I,G,Q)$, shown in Figure 11, is derived from Cuconasu et al. (2024), with slight modifications to better adapt to our setting.
@@
We follow the typical "retrieve-read" setting of RAG paradigm. For the retrieval module, we use Contriever-msmarco [^8] https://huggingface.co/facebook/contriever-msmarco, a BERT-based dense retriever, as the default retriever. It is finetuned on the MS MARCO dataset Bajaj et al. (2016) after unsupervised pretraining via contrastive learning Izacard et al. (2021). To optimize the efficiency of vector similarity searches, we employ the Faiss library Douze et al. (2024). For the read module, we deploy LLMs on NVIDIA A100 GPUs and accelerate inference with vllm [^9] https://github.com/vllm-project/vllm. We set the temperature to 0.1 to ensure stable outputs and strong reproducibility.
@@
Figure 11: Instruction II used for the QA task.
@@
Taxonomy Perturbations Known-Golden Known-Noise Unknown-Golden U-N
LR $R$ WR Org Acc LR $R$ WR Org Acc LR $R$ WR Org Acc $R$
Style Simple 7.79 83.04 9.18 66.03 67.42 1.70 95.80 2.50 4.12 4.92 8.43 82.88 8.69 51.42 51.68 99.45
Complex 6.00 85.60 8.40 68.43 1.91 96.59 1.50 3.71 6.71 84.86 8.43 53.13 99.57
[1pt/1pt] Source LLM-Generated 5.89 86.43 7.69 65.62 67.43 1.43 96.83 1.74 4.13 4.45 6.20 85.71 8.09 49.15 51.04 99.56
Self-Generated 6.55 85.01 8.44 67.52 1.55 96.37 2.09 4.67 6.52 86.36 7.12 49.74 99.57
[1pt/1pt] Logic Reverse 5.06 90.82 4.12 62.95 62.01 1.13 97.82 1.06 4.43 4.36 5.73 89.71 4.56 45.84 44.66 99.67
Random 3.91 93.16 2.93 61.97 0.86 98.31 0.83 4.40 4.21 91.67 4.12 45.74 99.72
LLM-Ranked 3.24 93.93 2.83 62.54 0.82 98.43 0.74 4.36 3.58 93.36 3.06 45.32 99.76
[1pt/1pt] Format JSON 7.01 88.25 4.74 63.91 61.64 1.70 97.25 1.05 3.87 3.21 5.92 89.63 4.45 49.35 47.88 99.61
HTML 11.85 84.46 3.69 55.75 2.70 96.90 0.40 1.56 9.33 86.78 3.90 43.92 99.61
YAML 5.26 89.94 4.80 63.45 1.26 97.41 1.33 3.94 4.79 90.80 4.41 48.97 99.67
Markdown 2.32 92.23 5.45 67.04 0.60 96.89 2.51 5.77 2.34 93.46 4.19 51.20 99.61
[1pt/1pt] Metadata Timestamp (pre) 2.08 95.81 2.11 55.77 55.80 0.28 99.42 0.29 1.58 1.59 2.54 95.56 1.90 43.31 42.66 99.95
Timestamp (post) 2.04 95.86 2.10 55.84 0.25 99.43 0.32 1.64 2.81 95.56 1.63 42.12 99.95
Datasource (wiki) 2.11 93.45 4.44 58.10 0.23 98.96 0.81 2.17 3.25 92.47 4.27 44.33 99.86
Datasource (twitter) 2.27 94.11 3.62 57.11 0.31 99.25 0.43 1.70 2.77 93.97 3.25 43.79 99.91
@@
Appendix F Details of SIG_Trivial Benchmark
We additionally construct a new benchmark, SIG_Trivial, following the SURE framework pipeline. This new benchmark uses a production-level retriever (Bing Search) to retrieve documents from the open web, with queries sources, retrieval algorithms, and data sources all differing from those used in our original SIG dataset. Specifically, we sample 1000 queries from TriviaQA and then sample 3 web documents (longer length than passages from wikepdeia) with golden answers from Bing Search for each query to serve as the grounding data. For each subcategory, 100 instances are randomly sampled to construct the new benchmark.
Appendix G Implementation of LLM-as-Judge
Specifically, we provide two demonstrations and prompt the LLM (Llama-3.1-70B-Instruct) to act as an evaluator. Before reaching a final judgment, the model is instructed to reason step by step using the chain-of-thought (COT) approach, encouraging a more thorough consideration of answer quality.
Appendix H Training Setups
Table 10 presents the detailed hyper-parameters for our training. We conduct our experiments on a server equipped with eight NVIDIA A100-SXM4-80GB GPUs.
Hyper-parameter Value
Learning Rate $2\times 10^{-5}$
Number of Epochs 2
Number of Devices 4
Per-device Batch Size 4
Gradient Accumulation Steps 2
Effective Batch Size 32
Optimizer AdamW
Learning Rate Scheduler cosine
Warmup Steps 100
Max Sequence Length 4096
Table 10: The hyper-parameters used for full parameter fine-tuning of models.